x + 2y = -1
3x - 2y =9
SOLVE FOR X AND Y

Found 3 solutions by stanbon, algebrahouse.com, mananth:
Answer by stanbon(75887)

x + 2y = -1
3x - 2y = 9
-----
Add and solve for "x":
4x = 8
x = 2
----
Solve for "y":
x + 2y = -1
2 + 2y = -1
2y = -3
y = -3/2
===============
Cheers,
Stan H.
================

Answer by algebrahouse.com(1659)

x + 2y = -1
3x - 2y = 9
------------
4x = 8 {added the two equations together}
x = 2 {divided both sides by 4}

x + 2y = -1 {top equation}
2 + 2y = -1 {substituted 2, in for x, into top equation}
2y = -3 {subtracted 2 from both sides}
y = -3/2 {divided both sides by 2}

x = 2 and y = -3/2
www.algebrahouse.com

Answer by mananth(16946)

x+2y=-1 .............1
3x-2y=9 .............2

multiply (1)by 1
Multiply (2) by 1
x+2y=-1
3x-2y=9

Add the two equations

4x=8

/4

x=2

plug value of x in (1)

x+2y=-1
2+2y=-1
2y=-1+-2
2y=-3
y=-1.5